OpenStudy (anonymous):

Can some1 plez help, (b+1/2) (b+1/5)

OpenStudy (anonymous):

\[(b+\frac{1}{2})(b+\frac{1}{5})\] \[=b^2+\frac{1}{5}b+\frac{1}{2}b+\frac{1}{2}\times \frac{1}{5}\]

OpenStudy (anonymous):

so will that be the simplied answer

OpenStudy (anonymous):

\[=b^2+\frac{7}{10}b+\frac{1}{10}\]